April 2015
Thu
23
The day has come.

Found a good home for all inhabitants -- fish, corals, and every last member of the CUC. Aquascape given to LFS. Water is drained; sand is out; Sump is out back drying. Tomorrow the remainder of the breakdown continues.

The end of a fun, but sometimes frustrating project.
Tue
21
Cleaned glass
Changed socks
Mon
20
Measured pH at 8.39 pH
Measured Temperature at 77.8 F
Measured ORP at 284 mV
Sun
19
Measured pH at 8.43 pH
Measured Temperature at 77.7 F
Measured ORP at 284 mV
Sat
18
Measured ORP at 287 mV
Measured pH at 8.39 pH
Measured Temperature at 77.7 F
Changed socks
Changed 1 Gallon of water
  0.6%
Brushed GHA
Cleaned probes
Replaced ROX Carbon
Replaced High-cap GFO
Fri
17
Measured ORP at 322 mV
Measured Temperature at 77.7 F
Measured pH at 8.33 pH